# AR-LIF: Adaptive Reset Leaky Integrate-and-Fire Neuron for Spiking Neural Networks

### v2

@@ 2.1 Leaky Integrate-and-Fire (LIF) Neuron
$u^{l}[t]=k_{\tau}u^{l}[t-1]+i^{l}[t],$ (1)
where $i^{l}[t]=W^{l}*s^{l-1}[t]$, $W^{l}$ is the synaptic weight, $t$ is time step, and $l$ denotes the $l$-thth layer of the network. $k_{\tau}\in(0,1)$ is the neuron membrane decay factor, which is a constant. $i^{l}[t]$ is the input of the neuron, $u^{l}[t]$ is the membrane potential of the neuron, $s^{l-1}[t]$ is the spike output of the neuron.
Fire: When the membrane potential after leaky integration satisfies $u^{l}[t]>V_{\mathrm{th}}$ , the output is 1; otherwise, the output is 0.
$s^{l}[t]=\Theta\left(u^{l}[t]-V_{\text{th}}\right),$ (2)
@@
$u^{l}[t]=\left\{\begin{array}[]{c}u^{l}[t]\odot\left(1-s^{l}[t]\right),\text{ Hard Reset }\\ u^{l}[t]-\rho s^{l}[t],\text{ Soft Reset }\end{array}\right$. (3)
@@ 2.3 Adaptive Reset Leaky Integrate-and-Fire Neuron
Adaptive decay of input accumulations: Consider introducing an additional memory variable $r^{l}[t]$ into the LIF neuron. $r^{l}[t]$ has an attenuation property, and the strength of its attenuation is associated with the input. The larger the input at the current time step, the larger the attenuated $r^{l}[t]$ will be, and vice versa. It can be described as:
@@
where $\alpha^{l}\in\mathbb{R}$ is a learnable parameter, $\sigma(\cdot)$ is the sigmoid function, $i^{l}[t]$ is the input at time step $t$, and $r^{l}[-1]=0$.
Spike feedback calculation of input accumulations: $r^{l}[t]$ is used to count the input penalty amount related to spikes and the input incentive amount without spikes. For inputs that generate spikes, they are accumulated normally; for inputs that do not generate spikes, negative feedback accumulation is performed. It can be described as:
@@
where $V_{\mathrm{th}}$ is the neuron threshold, which is usually set to 1. Here, we enable the threshold to undergo small-range adaptive adjustment; specifically, $V_{\text{th}}^{l}[t]=1+0.1\cdot\beta^{l}\cdot\operatorname{Tanh}(i^{l}[t])$, $\operatorname{Tanh}(\cdot)$ is the hyperbolic tangent function, $\beta^{l}\in[-1,1]$.
@@
$u^{l}[t]=h^{l}[t]-v_{\mathrm{r}}^{l}[t],$ (8)
where $u^{l}[t]$ is the membrane potential after spike firing, $h^{l}[t]$ is the integrated membrane potential, $v_{\mathrm{r}}^{l}[t]$ is the adaptive reset voltage.
@@
$\left\{\begin{array}[]{c}h^{l}[t]=k_{\tau}u^{l}[t-1]+i^{l}[t],\\ s^{l}[t]=\Theta\left(h^{l}[t]-V_{\text{th}}^{l}[t]\right),\\ r^{l}[t]=\left\{\begin{array}[]{c}\sigma\left(\alpha^{l}\cdot i^{l}[t]\right)\odot r^{l}[t-1],r^{l}[t-1]\geq 0,\\ \left(1-\sigma\left(\alpha^{l}\cdot i^{l}[t]\right)\right)\odot r^{l}[t-1],r^{l}[t-1]<0,\end{array}\right.\\ v_{\mathrm{r}}^{l}[t]=r^{l}[t]+\left(2s^{l}[t]-1\right)\odot\sigma\left(i^{l}[t]\right),\\ u^{l}[t]=h^{l}[t]-\left(v_{\mathrm{r}}^{l}[t]+V_{\text{th}}^{l}[t]\right),\end{array}\right$. (9)
@@ 3.1 Experimental Details
$\frac{\partial s^{l}[t]}{\partial h^{l}[t]}=\frac{1}{a}\cdot\operatorname{sign}\left(\left|v^{l}[t]-V_{\mathrm{th}}\right|<\frac{a}{2}\right),$ (12)
where $a$ is set to 1. When $V_{\text{th}}-0.5\leq v^{l}[t]\leq V_{\text{th}}+0.5$, $\frac{\partial s^{l}[t]}{\partial h^{l}[t]}=1$ ; otherwise, $\frac{\partial s^{l}[t]}{\partial h^{l}[t]}=0$.
@@ 3.2 Comparisons with Existing Methods
We evaluated the performance of AR-LIF neurons on multiple static and neuromorphic datasets, and the results of their comparison with SOTA methods are presented in Table 1.
@@
Table 2: Ablation study. ”Both” denotes the scenario where $\alpha=1$ and $V_{\text{th}}^{l}[t]=1$, while ”Learnable” denotes AR-LIF.
Dataset
4 Time Steps
Accuracy(%)
@@ 4 Time Steps
$\alpha=1$ $V_{\text{th}}^{l}[t]=1$ Both Learnable

### v3

@@ 2.1 Leaky Integrate-and-Fire (LIF) Neuron
$h^{l}[t]=k_{\tau}u^{l}[t-1]+i^{l}[t],$ (1)
where $i^{l}[t]=W^{l}*s^{l-1}[t]$, $W^{l}$ is the synaptic weight, $t$ is time step, and $l$ denotes the $l$-thth layer of the network. $k_{\tau}\in(0,1)$ is the neuron membrane decay factor, which is a constant. $i^{l}[t]$ is the input of the neuron, $u^{l}[t-1]$ is the membrane potential after spike firing, $h^{l}[t]$ is the integrated membrane potential, $s^{l-1}[t]$ is the spike output of the neuron.
Fire: When the membrane potential after leaky integration satisfies $h^{l}[t]>V_{\mathrm{th}}$ , the output is 1; otherwise, the output is 0.
$s^{l}[t]=\Theta\left(h^{l}[t]-V_{\text{th}}\right),$ (2)
@@
$u^{l}[t]=\left\{\begin{array}[]{c}h^{l}[t]\odot\left(1-s^{l}[t]\right),\text{ Hard Reset }\\ h^{l}[t]-\rho s^{l}[t],\text{ Soft Reset }\end{array}\right$. (3)
@@ 2.3 Adaptive Reset Leaky Integrate-and-Fire Neuron
Adaptive decay of input accumulations: Consider introducing an additional memory variable $r^{l}[t]$ into the LIF neuron. $r^{l}[t]$ has an decay property, and the strength of its decay is associated with the input. For negative values of $r^{l}[t]$, a complementary decay coefficient is adopted. It can be described as:
@@
where $R^{l}[t]$ denotes an intermediate quantity prior to the accumulation that yields $r^{l}[t]$, $\alpha^{l}\in\mathbb{R}$ is a learnable parameter, $\sigma(\cdot)$ is the sigmoid function, $i^{l}[t]$ is the input at time step $t$, and $r^{l}[-1]=0$.
Spike feedback calculation of input accumulations: $r^{l}[t]$ is used to count the input penalty amount related to spikes and the input incentive amount without spikes. For inputs that generate spikes, they are accumulated normally; for inputs that do not generate spikes, reverse accumulation is performed. It can be described as:
@@
where $V_{\mathrm{th}}$ is the neuron threshold, which is usually set to 1. We enable the threshold to undergo small-range adaptive adjustment; specifically, $V_{\text{th}}^{l}[t]=1+\beta^{l}\operatorname{Tanh}(i^{l}[t])$, $\operatorname{Tanh}(\cdot)$ is the hyperbolic tangent function, $\beta^{l}\in(-1,1)$.
@@
$u^{l}[t]=h^{l}[t]-v_{\mathrm{r}}^{l}[t]\cdot s^{l}[t],$ (8)
where $v_{\mathrm{r}}^{l}[t]$ is the adaptive reset voltage.
@@
$\left\{\begin{array}[]{c}h^{l}[t]=k_{\tau}u^{l}[t-1]+i^{l}[t],\\ s^{l}[t]=\Theta\left(h^{l}[t]-V_{\text{th}}^{l}[t]\right),\\ R^{l}[t]=\left\{\begin{array}[]{c}\sigma\left(\alpha^{l}i^{l}[t]\right)\cdot r^{l}[t-1],r^{l}[t-1]\geq 0,\\ \left(1-\sigma\left(\alpha^{l}i^{l}[t]\right)\right)\cdot r^{l}[t-1],r^{l}[t-1]<0,\end{array}\right.\\ r^{l}[t]=R^{l}[t]+(2s^{l}[t]-1)\cdot\sigma(i^{l}[t]),\\ v_{\mathrm{r}}^{l}[t]=V_{\mathrm{th}}^{l}[t]+\sigma(r^{l}[t]),\\ u^{l}[t]=h^{l}[t]-v_{\mathrm{r}}^{l}[t]\cdot s^{l}[t],\end{array}\right$. (9)
@@ 3.1 Experimental Details
$\frac{\partial s^{l}[t]}{\partial h^{l}[t]}=\frac{1}{a}\cdot\operatorname{sign}\left(\left|h^{l}[t]-V_{\mathrm{th}}\right|<\frac{a}{2}\right),$ (12)
where $a$ is set to 1. When $V_{\text{th}}-0.5\leq h^{l}[t]\leq V_{\text{th}}+0.5$, $\frac{\partial s^{l}[t]}{\partial h^{l}[t]}=1$ ; otherwise, $\frac{\partial s^{l}[t]}{\partial h^{l}[t]}=0$.
@@ 3.2 Comparisons with Existing Methods
We evaluated the performance of AR-LIF on multiple static and neuromorphic datasets, and the results of their comparison with SOTA methods are presented in Table 1.
@@
Table 2: Ablation study. ”Both” denotes the scenario where $\alpha^{l}=1$ and $V_{\text{th}}^{l}[t]=1$, while ”Learnable” denotes AR-LIF.
Dataset 4 Time Steps Accuracy(%)
@@
$\alpha^{l}=1$ $V_{\text{th}}^{l}[t]=1$ Both Learnable